Using the method of Laplace transforms, solve the following initial value problem:

y"-3y'+2y=6e^-t; y(0)=3,y'(0)=3
1
Expert's answer
2020-03-04T09:07:10-0500

Taking Laplace transform, we get;

"(s^2y^{-}(s)-sy(0)-y'(0))-3(sy^{-}(s)-y'(0))+2y^{-}(s)=L(6e^{-t})"

"(s^2y^{-}(s)-3s-3)-3(sy^{-}(s)-3)+2y^{-}(s)=6\/(s+1)"

"(s^2-3s+2)y^{-}(s)=3(s-2)+6\/(s+1)=(3s^2-3s)\/(s+1)"

"\implies y^{-}(s)=3s/(s+1)(s-2)"

Taking inverse Laplace transform, we get;

"y(x)=L^{-1}(3s\/(s+1)(s-2))"

"y(x)=L^{-1}(A\/(s+1)+B\/(s-2))"

"\implies 3s=A(s-2)+B(s+1)"

Equating the coefficients to solve for A,B we get;

"A+B=3;-2A+B=0"

Solving these equations, we get; "A=1;B=2"

"y(x)=L^{-1}((1\/(s+1)+2\/(s-2))"

"=L^{-1}(1\/(s-(-1)))+2L^{-1}(1\/(s-2))"

"=e^{-x}+2e^{2x}" ---(Answer)